# Positional Preservation Embedding for Multimodal Large Language Models

### v1

@@ Abstract
Multimodal large language models (MLLMs) have achieved strong performance on vision-language tasks, yet often suffer from inefficiencies due to redundant visual tokens. Existing token merging methods reduce sequence length but frequently disrupt spatial layouts and temporal continuity by disregarding positional relationships. In this work, we propose a novel encoding operator dubbed as Positional Preservation Embedding (PPE), which has the main hallmark of preservation of spatiotemporal structure during visual token compression. PPE explicitly introduces the disentangled encoding of 3D positions in the token dimension, enabling each compressed token to encapsulate different positions from multiple original tokens. Furthermore, we show that PPE can effectively support cascade clustering — a progressive token compression strategy that leads to better performance retention. PPE is a parameter-free and generic operator that can be seamlessly integrated into existing token merging methods without any adjustments. Applied to state-of-the-art token merging framework, PPE achieves consistent improvements of $2\%\sim 5\%$ across multiple vision-language benchmarks, including MMBench (general vision understanding), TextVQA (layout understanding) and VideoMME (temporal understanding). These results demonstrate that preserving positional cues is critical for efficient and effective MLLM reasoning.
@@ 1 Introduction
In the experiments, we apply PPE for tackling a variety of vision-language tasks, including general vision-language understanding on MMBench Liu et al. (2024), text-based visual reasoning on TextVQA Singh et al. (2019), temporal understanding on VideoMME Fu et al. (2024), etc. The reported performances consistently outstrip previous compression methods Dhouib et al. (2025); Jin et al. (2024) by significant margins after fine-tuning. We strongly believe that PPE can make inroads into domains of sparse token representing in MLLM where dense visual representation had previously reigned supreme.
@@ 2.2 Visual Token Merging
To improve the efficiency of MLLMs, recent works have explored visual token reduction strategies such as clustering and merging Bolya et al. (2023a); Xu et al. (2022); Ma et al. (2023); Zeng et al. (2022); Dhouib et al. (2025); Song et al. (2024); Jin et al. (2024). These methods significantly reduce the sequence length of image or video inputs, enabling faster inference and reduced memory consumption. Techniques include token matching Bolya et al. (2023a), hierarchical clustering Ma et al. (2023); Zeng et al. (2022), group-based representation Xu et al. (2022), and pruning-clustering hybrids Dhouib et al. (2025), with extensions to long-video understanding via sparse memory representations Song et al. (2024) and unified token compression across modalities Jin et al. (2024). Despite their effectiveness, these methods typically discard original positional information, which can disrupt spatial layouts and temporal continuity—limiting performance on fine-grained visual reasoning tasks. In contrast, our approach explicitly preserves spatiotemporal position cues throughout compression, maintaining structural fidelity while achieving substantial token reduction.
@@ 4.1 Experimental Setup
Evaluation Benchmarks. We use VideoMME Fu et al. (2024) as our primary benchmark for multimodal video analysis. To assess generalization, we also report results on NeXT-QA (multi-choice and open-ended) Xiao et al. (2021), SEED‑Bench‑Video Li et al. (2023a), and MVBench Li et al. (2024b). For image understanding, we evaluate on MMBench (CN/EN) Liu et al. (2024), SQA Iyyer et al. (2017), and TextVQA Singh et al. (2019).
@@ 4.2 Main Results
Table 1: The overall performance comparison across different benchmarks. The Dense model is a simply Qwen2.5-VL-3B model fine-tuned on our SFT datasets. The subscript S/ST indicates spatial-only and spatiotemporal compression, respectively.
@@
Image Average 82.09 75.90 81.15 (+5.25) - -
@@
Image Tasks. Table 1 illustrates the strong performance of PPE on image understanding benchmarks. Despite reducing visual tokens by 55%, PPE achieves competitive or superior performance across all tasks. It significantly outperforms Chat-UniVi in overall accuracy (81.15% vs. 75.9%) and on TextVQA (77.14% vs. 57.66%), while maintaining comparable results on other tasks. Compared to the full-token Dense baseline model, PPE shows only a minor performance drop in overall accuracy (81.15% vs. 82.09%), despite using less than half the visual tokens.
@@ 4.4 Ablation Study
In this section, we study the insight of PPE, its compatibility with other clustering-based compression methods, inference efficiency, and the effect of retaining different numbers of token ID positions to verify its ability to preserve positional information. Further studies on different reduction ratios A.1, aggregation stages A.2, backbone generalization A.3, and task generalization A.4, are provided in the Appendix A.
4.4.1 Insight into why PPE better preserves token semantics.
@@ 4.4.4 Effect of $K$ on Performance and Retained ID Ratio
The hyperparameter $K$ determines how many token ID positions are retained after merging, thus affecting spatiotemporal preservation. As shown in Table 5 and Figure 1(b), $K=1$ causes severe degradation due to insufficient positional information, while $K=24$ introduces redundancy and slightly underperforms $K=8$. The choice of $K=8$ achieves the best trade-off, aligning with the GCD of the 3D M-RoPE dimensions [16, 24, 24] to preserve essential signals without over-retention, and is adopted as our default. Moreover, Table 5 shows that larger $K$ values yield higher valid ID retention; for $K>1$, the retention rate exceeds the visual token retention ratio (45%), confirming that PPE captures positional information beyond token counts.
@@ A.3 Analysis of Backbone Generalization
PPE (Ours) 56.40 42.10 35.70 44.74 96%
A.4 Analysis of Task Generalization
@@
A.5 Qualitative Cases on Image Captioning
@@
A.6 Attention Visualization Analysis
@@
A.7 Limitations and Failure Cases
@@
As shown in Tables 1 and 6, PPE underperforms Dense in certain cases, highlighting the trade-off between compression and fidelity. We visualize representative failure cases to illustrate PPE’s limitations.

### v2

# PPE: Positional Preservation Embedding for Token Compression in Multimodal Large Language Models

@@ Abstract
Multimodal large language models (MLLMs) have achieved strong performance on vision-language tasks, yet often suffer from inefficiencies due to redundant visual tokens. Existing token merging methods reduce sequence length but frequently disrupt spatial layouts and temporal continuity by disregarding positional relationships. In this work, we propose a novel encoding operator dubbed as Positional Preservation Embedding (PPE), which has the main hallmark of preservation of spatiotemporal structure during visual token compression. PPE explicitly introduces the disentangled encoding of 3D positions in the token dimension, enabling each compressed token to encapsulate different positions from multiple original tokens. Furthermore, we show that PPE can effectively support cascade clustering — a progressive token compression strategy that leads to better performance retention. PPE is a parameter-free and generic operator that can be seamlessly integrated into existing token merging methods without any adjustments. Applied to state-of-the-art token merging framework, PPE achieves consistent improvements of $2\%\sim 5\%$ across multiple vision-language benchmarks, including MMBench (general vision understanding), TextVQA (layout understanding) and VideoMME (temporal understanding). These results demonstrate that preserving positional cues is critical for efficient and effective MLLM reasoning. Our code is available at https://github.com/MouxiaoHuang/PPE.
@@ 1 Introduction
In the experiments, we apply PPE for tackling a variety of vision-language tasks, including general vision-language understanding on MMBench Liu et al. (2024a), text-based visual reasoning on TextVQA Singh et al. (2019), temporal understanding on VideoMME Fu et al. (2024), etc. The reported performances consistently outstrip previous compression methods Dhouib et al. (2025); Jin et al. (2024) by significant margins after fine-tuning. We strongly believe that PPE can make inroads into domains of sparse token representing in MLLM where dense visual representation had previously reigned supreme.
@@ 2.2 Visual Token Merging
To improve the efficiency of MLLMs, recent works have explored visual token reduction strategies Kong et al. (2025) such as clustering and merging Bolya et al. (2023a); Xu et al. (2022); Ma et al. (2023); Zeng et al. (2022); Dhouib et al. (2025); Song et al. (2024); Jin et al. (2024); Yang et al. (2025). These methods significantly reduce the sequence length of image or video inputs, enabling faster inference and reduced memory consumption. Techniques include token matching Bolya et al. (2023a), hierarchical clustering Ma et al. (2023); Zeng et al. (2022), group-based representation Xu et al. (2022), and pruning-clustering hybrids Dhouib et al. (2025), with extensions to long-video understanding via sparse memory representations Song et al. (2024) and unified token compression across modalities Jin et al. (2024). Despite their effectiveness, these methods typically discard original positional information, which can disrupt spatial layouts and temporal continuity—limiting performance on fine-grained visual reasoning tasks. In contrast, our approach explicitly preserves spatiotemporal position cues throughout compression, maintaining structural fidelity while achieving substantial token reduction.
@@ 4.1 Experimental Setup
Evaluation Benchmarks. We use VideoMME Fu et al. (2024) as our primary benchmark for multimodal video analysis. To assess generalization, we also report results on NeXT-QA (multi-choice and open-ended) Xiao et al. (2021), SEED‑Bench‑Video Li et al. (2023a), and MVBench Li et al. (2024b). For image understanding, we evaluate on MMBench (CN/EN) Liu et al. (2024a) and SQA Iyyer et al. (2017) to assess general visual comprehension. To further examine text-rich layout and document understanding, we include TextVQA Singh et al. (2019), DocVQA Mathew et al. (2021), OCRBench Liu et al. (2024b), and ChartQA Masry et al. (2022). All reported metrics follow the evaluation criteria of their respective benchmarks, and unless otherwise noted, higher values indicate better performance.
@@ 4.2 Main Results
Table 1: The overall performance comparison across different benchmarks. The Dense model is a simply Qwen2.5-VL-3B-Instruct model fine-tuned on our SFT datasets. The subscript S/ST indicates spatial-only and spatiotemporal compression, respectively. Bold denotes the best performance among compressed variants, and underline denotes the best overall including the Dense baseline. The image average score is calculated without OCRBench.
@@
DocVQA 89.44 52.48 76.79 (+24.31) - -
OCRBench 761 535 691 (+156) - -
ChartQA 79.96 49.60 74.52 (+24.92) - -
Image Average 82.96 67.61 79.32 (+11.71) - -
@@
Image Tasks. Table 1 illustrates the strong performance of PPE on image understanding benchmarks. Despite reducing visual tokens by 55%, PPE achieves competitive or superior performance across all tasks. It significantly outperforms Chat-UniVi in overall accuracy (79.32% vs. 67.61%) and on TextVQA (77.14% vs. 57.66%), while maintaining comparable results on other tasks. Compared to the full-token Dense baseline model, PPE shows only a minor acceptable performance drop in overall accuracy (82.96% vs. 79.32%), despite using less than half the visual tokens.
@@ 4.4 Ablation Study
In this section, we study the insight of PPE, its compatibility with other clustering-based compression methods, inference efficiency, and the effect of retaining different numbers of token ID positions to verify its ability to preserve positional information. Further studies on different reduction ratios A.1, aggregation stages A.2, model size generalization A.4, backbone generalization A.3, task generalization A.5, etc, are provided in the Appendix A.
4.4.1 Insight into why PPE better preserves token semantics
@@ 4.4.2 Integration with Other Clustering-Based Methods
To further demonstrate PPE’s generality across different compression frameworks, we conduct extensive experiments on VisionZip Yang et al. (2025), which combines pruning and merging and supports both training-free and SFT settings. We simply integrate PPE into VisionZip on Qwen2.5-VL-3B-Instruct following the official instructions (dominant_ratiodominant\_ratio and contextual_ratiocontextual\_ratio are set $0.40$ and $0.05$ respectively to achieve 55% reduction ratio), randomly sampling 1/10 of our curated dataset for the SFT setting and only fine-tuning the projector for one epoch. The results are shown in Table 5. In both settings, PPE delivers improvements in most cases, especially on layout- and OCR-sensitive benchmarks, demonstrating that PPE retains more spatial information under the same reduction ratio. Moreover, when training is allowed, PPE achieves larger gains. Our method is decoupled from model size., extensive experiments are provided in Appendix A.4.
Table 5: Extensive comparisons under both training-free and SFT using Qwen2.5-VL-3B-Instruct.
Setting Benchmarks Dense Chat-UniVi Chat-UniVi+PPE VisionZip VisionZip+PPE
Training-Free MMBench (EN) 79.10 81.50 82.28 (+0.78) 83.48 83.18 (-0.30)
MMBench (CN) 78.10 80.06 81.43 (+1.37) 81.75 81.64 (-0.11)
TextVQA 79.30 37.60 73.78 (+36.18) 79.00 79.62 (+0.62)
DocVQA 93.90 19.58 66.16 (+46.58) 83.98 85.63 (+1.65)
OCRBench 797 307 598 (+291) 713 725 (+12)
ChartQA 84.00 18.72 67.08 (+48.52) 79.72 80.72 (+1.00)
SFT MMBench (EN) 85.89 84.92 84.73 (-0.19) 83.99 83.78 (-0.21)
MMBench (CN) 86.07 83.71 84.87 (+1.16) 82.12 82.63 (+0.15)
TextVQA 79.50 57.66 77.14 (+19.48) 80.02 82.06 (+2.04)
DocVQA 89.44 52.48 76.79 (+24.31) 84.84 90.52 (+5.68)
OCRBench 761 535 691 (+156) 711 780 (+69)
ChartQA 79.96 49.60 74.52 (+24.92) 80.20 82.36(+2.16)
Reduction Ratio 0% 55% 55% 55% 55%
4.4.3 SFT or Training-Free
PPE is a plug-and-play, parameter-free method. Whether SFT is required depends on the underlying compression method, not PPE itself: 1) Chat-UniVi is a merging-based method that requires SFT, and performs bad in training-free setting especially on layout-sensitive benchmarks. 2) PACT and ToMe are fully training-free, and PPE is directly applied to the official pre-trained models without any additional training. 3) VisionZip supports both training-free and SFT usage, and PPE improves performance in both settings without modifying their token selection or aggregation mechanisms. This supports our key claim that PPE can enhance existing token compression pipelines in a plug-and-play manner and can further improve RoPE allocation when training is allowed.
@@ 4.4.5 Effect of $K$ on Performance and Retained ID Ratio
The hyperparameter $K$ determines how many token ID positions are retained after merging, thus affecting spatiotemporal preservation. As shown in Table 6 and Figure 1(b), $K=1$ causes severe degradation due to insufficient positional information, while $K=24$ introduces redundancy and slightly underperforms $K=8$. The choice of $K=8$ achieves the best trade-off, aligning with the GCD of the 3D M-RoPE dimensions [16, 24, 24] to preserve essential signals without over-retention, and is adopted as our default. Moreover, Table 6 shows that larger $K$ values yield higher valid ID retention which refers to the ratio of distinct positional embeddings preserved after compression; for $K>1$, the retention rate exceeds the visual token retention ratio (45%), confirming that PPE captures positional information beyond token counts. Further visualization analyses for different choices of $K$ are presented in Appendix A.10.
@@ A.3 Analysis of Backbone Generalization
PPE (Ours) 56.40 (+3.40) 42.10 (+7.70) 35.70 (-0.10) 44.74 (+1.67) 96%
A.4 Analysis of Model Size
Our method is decoupled from model size. To further demonstrate that PPE scales with model size, we additionally evaluate PPE on Qwen2.5-VL-7B-Instruct, results are shown in Table 10, showing that PPE provides improvements regardless of model capacity. For the SFT setting, we trained it on only 1/5 of the data used for the 3B model due to time and resource constraints.
We also note that in the training-free setting of VisionZip, PPE shows a slight performance drop on ChartQA for the 7B model, while improving the 3B model. This may be because the stronger 7B baseline already leverages fine-grained local patterns for such task, and without SFT, the additional positional details introduced by PPE can slightly perturb the attention distribution. After SFT, PPE consistently brings positive gains. For instance, the slight drop on VisionZip+PPE for ChartQA becomes an improvement, indicating that PPE benefits more when SFT is applied.
Table 10: Ablation study on model size using Qwen2.5-VL-7B-Instruct.
Setting Benchmarks Dense Chat-UniVi Chat-UniVi+PPE VisionZip VisionZip+PPE
Training-Free MMBench (EN) 83.50 83.23 83.58 (+0.35) 85.24 85.19 (-0.05)
MMBench (CN) 83.40 80.18 82.35 (+2.17) 83.62 83.92 (+0.30)
TextVQA 84.90 35.82 63.44 (+27.42) 84.64 85.02 (+0.40)
DocVQA 95.70 27.06 66.42 (+39.36) 87.30 88.52 (+1.22)
OCRBench 864 479 577 (+98) 780 794 (+14)
ChartQA 87.30 19.92 46.72 (+29.90) 60.56 59.88 (-0.68)
SFT MMBench (EN) 86.90 86.23 86.26 (+0.03) 86.14 87.24 (+0.16)
MMBench (CN) 85.35 84.25 84.85 (+0.60) 85.12 85.15 (0.03)
TextVQA 87.20 54.92 82.46 (+27.54) 87.08 87.24 (+0.16)
DocVQA 92.97 50.01 85.84 (+35.83) 93.16 93.57 (+0.41)
OCRBench 826 584 764 (+180) 828 828 (+0)
ChartQA 86.32 43.96 78.88 (+34.92) 85.40 86.08 (+0.68)
Reduction Ratio 0% 55% 55% 55% 55%
A.5 Analysis of Task Generalization
@@
A.6 Qualitative Cases on Image Captioning
@@
A.7 Attention Visualization Analysis
@@
A.8 Limitations and Failure Cases
@@
As shown in Tables 1 and 7, PPE underperforms Dense in certain cases, highlighting the trade-off between compression and fidelity. We visualize representative failure cases to illustrate PPE’s limitations.
@@
A.9 PPE Gains across Different Image Resolutions
To evaluate the effect of PPE under varying image resolutions, we calculated the average resolution of TextVQA, MMBench, DocVQA, OCRBench, and ChartQA, based on the resolution statistics of their samples. As illustrated in Figure 7, the horizontal axis denotes the average image resolution for each dataset, while the left vertical axis presents the performance difference ($\Delta$). This metric reflects both the improvements introduced by PPE over Chat-UniVi and its ability to recover performance relative to the dense model. For OCRBench, whose raw score range is not from 0 to 100, scores are normalized by a factor of 10 to ensure consistent visualization. The right vertical axis represents the number of image samples in each dataset.
From the figure, we observe that PPE consistently provides notable performance gains across a wide spectrum of image resolutions: from $385^{2}$ to $1924^{2}$. These results demonstrate that PPE imposes no limitations on input resolution and remains robust and effective across images of different scales.
Figure 7: PPE obtains consistent improvements across different resolutions.
A.10 Visualization of the Effect of $K$
To better understand how different values of $K$ influence attention behavior, we conduct a visual analysis and support it with quantitative indicators such as entropy and variance. Following the analytical framework introduced in Section 4.4.1 and Figure 3, we compare different choices of $K$ directly on the images.
As shown in Figure 8, when $K=1$, entropy is slightly higher and variance is lower, indicating that the model’s attention becomes more dispersed compared with $K=16$ or $K=32$. The visualizations further reveal that at $K=1$, the model mainly focuses on the first half of the brand text (with brighter activations in red). Meanwhile, tokens on the camera backplate also become more activated, showing that attention is distracted by background regions and fails to capture the full brand string, unlike the cases with $K=16$ or $K=32$. For images, we set $K=32$ based on the GCD of M-RoPE configuration $[32,32]$, which maximizes the preservation of positional information per merged token. The same rationale applies to video settings, $K=8$ of $[16,24,24]$.
M-RoPE splits embedding dimensions into multiple periodic components (temporal, vertical, horizontal), assigning each axis contiguous frequency bands. Existing analysis (e.g., Figure 3 and 4 in Reference Huang et al. (2025)) shows that uneven allocation can lead to some axes occupying only high- or low-frequency channels, impairing long-range and spatial modeling. Choosing K based on the GCD ensures that each merged token preserves multiple positional anchors evenly across the frequency spectrum, preventing collapse into only high- or low-frequency bands. In short, $K$ is grounded in M-RoPE’s frequency structure, providing theoretical justification for PPE’s ability to preserve positional information after token merging.
In short, refer to the analysis of positional encoding in VLMs Huang et al. (2025), choosing an appropriate $K$ ensures that positional indices are distributed evenly across frequency bands; otherwise, some indices may collapse into only high- or low-frequency ranges, degrading positional fidelity.
(a) Entropy comparison ↓
(b) Variance comparison ↑
(c) PPE (K=1)
(d) PPE (K=16)
(e) PPE (K=32)
Figure 8: Attention statistics and visualizations of samples from TextVQA. (a–b) Quantitative comparison of entropy and variance. (c–e) Qualitative attention score visualizations of case 34602.